If the cost was $20 in 1998 and $40 in 2008, then what is the average rate of change of the cost for that time period?
Answer by Boreal(15235)

It doubles in 10 years
p=Poe^rt
p/Po=2
2=e^rt
ln of both sides, t=10
0,0693=10r
r=6.93%. It is an interest problem with continuous compounding,